OpenStudy (anonymous):

sara had 7/8 pound of turkey she used 2/3 for sandwiches how many pounds of turkey does sara have left

OpenStudy (anonymous):

7/8 - 2/3 (common denominator is 24) 21/24 - 16/24 = 5/24 check.. 5/24 + 2/3 = 7/8 5/24 + 16/24 = 21/24 21/24 = 21/24 (correct) she has 5/24 pounds left